Larry Mitchell invested part of his $ 35,000

advance at 7%

annual simple interest and the rest at 2%

annual simple interest. If his total yearly interest from both accounts was $900​, find the amount invested at each rate.

Answer by addingup(3677)

So he invested some of 35,000 at 7% and the rest at 2%. The rest means what's left over from the money he invested at 7%, correct?
Then, if he invested "some", let's call it x, at 7%:
0.07x
And what's left over would be 35,000-x, correct?
:
0.07x+0.02(35000-x) = 900 apply the Distributive Law of multiplication, distribute the 0.02:
0.07x+700-0.02x = 900
0.05x = 200
x = 4,000 this is how much he invested at 7%
35,000-4,000 = 31,000 this is how much he invested at 2%
------------------------------------
Check:
4,000*0.07 = 280
+
31,000*0.02 = 620
. . . . . . -------
Total int. = 900 Correct
:
Happy learning,
John